OpenStudy (anonymous):

given cos theta=-1/5 , pi

OpenStudy (amistre64):

use the pythag identitiy: cos^2 + sin^2 = 1

OpenStudy (amistre64):

and since theta is in the 3rd quadrant, the sign of sin will be negative

OpenStudy (anonymous):

what is theta though?

OpenStudy (anonymous):

sin theta...

OpenStudy (amistre64):

doesnt matter, you already know that value of cos regardless of theta

OpenStudy (amistre64):

if we arrange the identity to solve for sin we get: cos^2 + sin^2 = 1 sin^2 = 1 - cos^2 sin = sqrt(1-cos^2) and since you already know cos, plug it in to determine sin

OpenStudy (amistre64):

if the mathing method is a bit to abstract, we also have the option of simply drawing a circle and applying the trig definitions to it

OpenStudy (anonymous):

sounds good.

OpenStudy (amistre64):



OpenStudy (amistre64):

cos = nextto/hyp so we apply the values given to the triangle about t sin = over/hyp ; in this case sqrt(10-1)/5 and since we are confined to an area in the region where sin is negative; just make sure the end result has a negative sign attached to it

OpenStudy (amistre64):

umm, 5^2 = 25 ... not 10 :)

OpenStudy (amistre64):

sin = sqrt(25-1)/5 would definnatly be a better value